OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/fourthx + 2. y = 4x - 1 y = 4x - 3 y = 1/fourthx - 3 y = 1/fourthx - 1

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

when 32 lines are parallel to each other their slope are the same...so the slope of this line would be 1/4 too.

OpenStudy (anonymous):

*2 lines not 32 lines XD

OpenStudy (anonymous):

last one..D..

OpenStudy (anonymous):

and u have slope and u only want y-intercept...do u have any idea how to solve it?

OpenStudy (anonymous):

i have no idea how to do this lol but thanks

OpenStudy (anonymous):

since parallel lines must have the same slope we know the slope will be 1/4. now plug everything into the equation y-y1=m(x-x1) y-(-2)=1/4(x-(-4)) y+2=1/4(x+4) y+2=1/4x+1 y=1/4x-1 D is the correct answer

OpenStudy (anonymous):

oh..see step by step...1st find the slope of given eq .. by comparing it with y=mx+c here m=slope.

OpenStudy (anonymous):

is your picture Stephen Hawking @PFEH.1999

OpenStudy (anonymous):

well i'll give you some :D .... now the equation is y = 1/4x + b which b= y-intercept...you have the point (-4,-2) so u can say -2 = (1/4)(-4) + b now solve for b

OpenStudy (anonymous):

Yes ... @stupid12